OpenStudy (babynini):

Find the exact value of sin(pi/12) using the Addition/Sub formulas

OpenStudy (babynini):

@Astrophysics

OpenStudy (babynini):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sin\left(\frac{\pi}{12}\right) = \sin\left(\frac{\pi}{3} - \frac{\pi}{4}\right)\] then use the identity sin(x-y) = sin(x)*cos(y) - cos(x)*sin(y)

jimthompson5910 (jim_thompson5910):

alternatively, you can use the fact that \[\Large \sin\left(\frac{\pi}{12}\right) = \sin\left(\frac{1}{2}*\frac{\pi}{6}\right)\] and use the half angle identity for sine. But it looks like they want you to use the first method

OpenStudy (babynini):

ooh thanks! Yes, I got the 1/2 * pi/6 but couldn't figure that out with the addition/subtraction and had forgotten about the half angle identity. Thanks! :)

OpenStudy (babynini):

(sorry for taking long to respond, I was having dinner and forgot)